### v1

# AdaR E T A K E : Adaptive Redundancy Reduction to Perceive Longer for Video-language Understanding

Abstract
Multimodal Large Language Models (MLLMs) have revolutionized video understanding, yet are still limited by context length when processing long videos. Recent methods compress videos by leveraging visual redundancy uniformly, yielding promising results. Nevertheless, our quantitative analysis shows that redundancy varies significantly across time and model layers, necessitating a more flexible compression strategy. We propose AdaR E T A K E , a training-free method that flexibly reduces visual redundancy by allocating compression ratios among time and layers with theoretical guarantees. Integrated into state-of-the-art MLLMs, AdaR E T A K E improves processing capacity from 256 to 2048 frames while preserving critical information. Experiments on VideoMME, MLVU, LongVideoBench, and LVBench datasets demonstrate that AdaR E T A K E outperforms existing methods by 2.3% and 2.8% for 7B and 72B models, respectively, with even greater improvements of 5.9% and 6.0% on the longest LVBench. Our code is available at https://github.com/SCZwangxiao/video-FlexReduc.git.
AdaR E T A K E : Adaptive Redundancy Reduction to Perceive Longer for Video-language Understanding
Xiao Wang1111Equal contribution. [^2] Work done during an internship at Huawei.  Qingyi Si2111Equal contribution.  Jianlong Wu1333Corresponding author.  Shiyu Zhu3  Li Cao2  Liqiang Nie1333Corresponding author. 1Harbin Institute of Technology, Shenzhen 2Huawei Technologies Co., Ltd. 3Shandong University scz.wangxiao@gmail.com, siqingyi@huawei.com, wujianlong@hit.edu.cn xyzcaoli@outlook.com, nieliqiang@gmail.com
1 Introduction
In pursuit of general intelligence, Multimodal Large Language Models (MLLMs) Li et al. (2024b); Lin et al. (2024) have revolutionized video understanding. However, current MLLMs require hundreds of tokens to represent a single image Wang et al. (2024a); Li et al. (2024a), limiting video lengths to less than 10 minutes Shen et al. (2024).
Efforts to extend MLLMs’ capabilities for long videos include: agent systems Zhang et al. (2024a) which retrieve and interpret pre-segmented videos but remain constrained by single-model abilities. Techniques like length extrapolation Zhang et al. (2024c) and multi-modal sequence parallelism Xue et al. (2024) enhance usable video context length but introduce more visual redundancy. Rather than extending context length, compression-based methods reduce video tokens into shorter sequences by leveraging visual redundancy Bolya et al. (2022). Many approaches He et al. (2024a); Fei et al. (2024) train Q-Former Li et al. (2023a) to condense videos guided by language or learnable query tokens. Recent advancements Shen et al. (2024); Wang et al. (2024c) integrate compression into MLLM prefilling, yielding promising results.
Figure 1: AdaR
E
T
A
K
E
enables MLLM to perceive longer with fixed context length for video-language understanding.
In this work, we push the boundaries of compression-based methods in two key ways: first, by optimizing the compression algorithm with insights from quantitative analysis; and second, by scaling the number of frames processed to capture more information from the video.
To dive deeper into compression-based methods, we quantitatively analyze visual redundancy by examining the distribution of influential tokens (more likely to be preserved during compression) during MLLM inference, revealing significant variations across video timestamps and LLM layers. These findings show that previous methods with fixed compression ratios fail to capture the dynamic nature of visual redundancy, underscoring the need for a more flexible compression strategy. In light of this, we propose AdaR E T A K E , a training-free adaptive video redundancy reduction method. It features two modules: temporal-adaptive allocation, which adjusts compression ratios for video sequence features over time, and layer-adaptive allocation, which manages KV cache states across layers. For temporal allocation, we divide a long video into chunks and allocate a compression ratio for each chunk based on the similarity scores between adjacent frames. For layer allocation, we adjust compression ratios across layers based on video-prompt attention scores. Our theoretical analysis demonstrates that this approach reduces the upper bound of $L_{1}$ compression loss compared to uniform allocation. The combination of the above allocation determines a specific compression ratio for each chunk in each LLM layer. Finally, we apply chunked prefilling for all chunks and the prompt. During this process, the KV caches of each chunk are compressed iteratively based on the accumulated attention scores. AdaR E T A K E compresses long videos into shorter sequences, allowing to perceive more informative frames within a fixed GPU memory budget, thereby enhancing long video understanding.
We integrate AdaR E T A K E into cutting-edge MLLMs like QWen2-VL Wang et al. (2024a) and LLaVA-Video Zhang et al. (2024e), and conduct extensive experiments across various video understanding benchmarks, including VideoMME Fu et al. (2024), MLVU Zhou et al. (2024), LongVideoBench Wu et al. (2024), and LVBench Wang et al. (2024b). The results show that AdaR E T A K E significantly outperforms existing methods, achieving an average improvement of 2.3% and 2.8% across datasets for 7B and 72B models, respectively. On LVBench, the dataset with the longest average video length, the gains are even more pronounced, with improvements of 5.9% and 6.0%, respectively. Additionally, the results on needle QA and temporal grounding tasks further demonstrate that our approach effectively preserves fine-grained temporal grounding capabilities. The ablation study validates the effectiveness of our temporal and layer-adaptive budget allocation methods. Through comparison with other compression approaches, it further demonstrates the superiority of our method. In summary, our contributions are threefold:
  • We identify uneven visual redundancy across time and MLLM layers and develop AdaR E T A K E to adaptively reduce it, expanding MLLM capacity from 256 to 2048 frames for long video understanding.
  • We design temporal- and layer-adaptive allocation modules to allocate compression ratios across time and MLLM layers, respectively. Theoretical analysis demonstrates that the layer-wise allocation effectively minimizes the upper bound of compression loss.
  • Our approach achieves state-of-the-art performance, surpassing existing MLLMs by an average of 2.3% and 2.8% across 4 datasets for 7B and 72B models, respectively.
2 Related Work
2.1 MLLM for Long Videos
Most existing multi-modal large language models struggle with extreme token lengths when applied directly to long videos. A commonly used and computationally manageable context length for multimodal training is 8k Shen et al. (2024), which restricts video processing to a few minutes.
Early attempts developed video agent systems Zhang et al. (2024a); Wang et al. (2024d); Luo et al. (2024) that segment videos into shorter clips and use MLLMs with open-source tools for retrieval, aggregation, and interpretation. However, a single model’s capabilities remain limited, reducing overall effectiveness. Length extrapolation methods Zhang et al. (2024c); Shang et al. (2024); Wei and Chen (2024) extend context windows beyond training lengths, but GPU memory still limits context size. To address this, Xue et al. introduced LongVILA, a multi-modal sequence parallelism system that distributes computation across GPUs, but this adds communication overhead Li et al. (2023b), affecting efficiency. In contrast, compression-based methods condense video tokens into shorter sequences. Approaches He et al. (2024a); Fei et al. (2024); Cheng et al. (2024); Zeng et al. (2024a); Man et al. (2024); Han et al. (2024) use Q-Former Li et al. (2023a) for token compression, reducing redundancy by leveraging language or query tokens. However, Q-Former, trained from scratch, lacks the world knowledge embedded in LLMs, making these methods suboptimal. Recent advances Shu et al. (2024); Shen et al. (2024); Liu et al. (2024); Wang et al. (2024c) integrate compression within the LLM, achieving promising results.
2.2 Token Compression for MLLMs
Token compression methods for LLMs Xiao et al. (2024); Zhang et al. (2023); Feng et al. (2024) reduce sequence length by evicting less important tokens, typically with some performance loss. Given the higher redundancy in visual tokens compared to language tokens Bolya et al. (2022), these methods have been extended to MLLMs Chen et al. (2024); Ye et al. (2024); He et al. (2024b); Zhu et al. (2024). Advancements include merging evicted tokens to reduce information loss Wan et al. (2024); Zhang et al. (2024d) and analyzing redundancy across layers Xing et al. (2024); Tu et al. (2024). However, unlike our adaptive allocation approach, these methods fail to exploit temporal redundancy and allocate compression ratios either monotonically Xing et al. (2024) or via heuristics Tu et al. (2024), resulting in suboptimal performance.
In this paper, we advance token compression methods for MLLMs by adaptively adjusting the compression ratio across timestamps and layers to reduce redundancy more effectively.
3 Preliminary Analysis
In this section, we provide a quantitative analysis of the visual redundancy with MLLM for long video understanding. Intuitively, redundancy varies across dimensions: at the frame level, static scenes are more redundant than dynamic ones, and at the model level, deeper layers focus on more abstract features, leading to different attention patterns. To quantify this, we measure redundancy through the ratio of heavy-hitters Zhang et al. (2023), a set of influential tokens essential for generation. By identifying these across dimensions, we validate the varying redundancy levels, providing a strong motivation for our approach to achieve more flexible and efficient compression.
Figure 2: Illustrating example of a heavy hitter. We adopt the heavy hitter ratio to measure the redundancy
Heavy-hitter ratio to measure redundancy. Denote the number of attention heads as $h$, length of prompt and video tokens as $L_{t}$ and $L_{v}$, respectively, and the attention scores of them in layer $l$ is $\textbf{A}^{(l)}\in\mathbb{R}^{h\times L_{t}\times L_{v}}$. We first calculate the prompt-accumulated head-average attention scores to measure the influence of each video token during generation $\mathbf{a}\in\mathbb{R}^{L_{v}}$:
$\mathbf{a}^{(l)}=\sum_{j=1}^{L_{t}}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}^{(l)}[i,j]}.$ (1)
We then calculate the heavy-hitter ratio $\lambda^{(l)}\in\mathbb{R}$:
$\lambda^{(l)}=\frac{1}{L_{v}}\sum_{i=1}^{L_{v}}{\mathbb{1}{\left(\mathbf{a}^{(l)}[i]>p\max{\left\{\mathbf{a}^{(l)}\right\}}\right)}},$ (2)
where $\mathbb{1}(\cdot)\in\{0,1\}$ is the indicator function and $p=0.01$ is a heuristic constant.A video token is considered important (called a heavy-hitter) if its accumulated attention $\mathbf{a}^{(l)}[i]$ exceeds $p$ times the maximum attention value in $\mathbf{a}^{(l)}$.
Figure 3: Heavy-hitter ratio among timestamps, showing the unevenly distributed temporal redundancy. The horizontal shaded bars indicate timestamps where the video has ended.
Figure 4: Heavy-hitter ratio among layers, showing the unevenly distributed redundancy among LLM layers.
Redundancy among video timestamps. To explore the distribution of redundancy over time, we first split the video tokens into chunks of 10 seconds, and denote the heavy hitter ratio chunk $t$ as $\lambda^{(t,l)}$. We randomly sampled 64 videos from VideoMME Fu et al. (2024) and plotted the layer-averaged heavy hitter ratio $\sum_{k}\lambda^{(t,l)}$ across different chunks as a heatmap in Figure 3. The temporal redundancy is unevenly distributed, with the heavy-hitter ratio varying up to 3x within a video, as highlighted by the red circle in Figure 2.
Redundancy among LLM layers. To investigate the distribution of redundancy across LLM layers in MLLM, we utilized all videos from VideoMME Fu et al. (2024) and plotted heavy hitter ratio $\sum_{k}\lambda^{(t,l)}$ across different layers as a boxplot in Figure 4. The redundancy is unevenly distributed among the LLM layers. Generally, the heavy hitter ratio is lower in the deeper layers, but significant fluctuations are observed, with local minima at layers 2, 14, and 21, and maxima at layers 7 and 18. This indicates that token compression methods that monotonically assign higher compression ratios to deeper layers, such as PyramidDrop Xing et al. (2024), are suboptimal for video understanding.
To maximize the use of informative frames within a fixed GPU memory budget, we must design a video compression algorithm that adaptively adjusts the compression ratio across different timestamps and LLM layers.
4 Methods
Figure 5: Illustration of AdaR
E
T
A
K
E
4.1 Overview
The architecture of AdaR E T A K E is shown in Figure 5. To flexibly reduce redundancy across timestamps, we divide video sequences into equal chunks and the Temporal-adaptive Allocation module dynamically applies distinct compression ratios to each chunk. For redundancy across layers, the Layer-adaptive Allocation module assigns varying compression ratios to LLM layers. Finally, the Token Compression module compresses the KV cache after each chunk’s prefilling based on the compression ratios determined by the previous modules, reducing the video sequence length in an MLLM. The general pipeline and these three modules are detailed below.
4.2 General Pipeline
Denote $T$ number of frames, $N$ number of tokens in each frame, $\tau$ number of frames in a chunk (can divide $T$), $L$ prompt length, and $C_{max}$ is a refined maximal context length.
Given raw frames and a text prompt as input, the visual encoder and projection layer derive video features $\mathbf{M}\in\mathbb{R}^{T\times N\times d}$, and the word embedding layer derives prompt features $\mathbf{P}\in\mathbb{R}^{L\times d}$. We split visual features into chunks of $\tau$ frames:
$\mathcal{M}=\left[\mathbf{M}_{1},\mathbf{M}_{2},...,\mathbf{M}_{T/\tau}\right],\mathbf{M}_{i}\in\mathbb{R}^{\tau\times N\times d}.$ (3)
The temporal-adaptive allocation module will produce a compression ratio (length after compression/original length) for each chunk based on the number of tokens in $\mathcal{M}$ and $C_{max}$:
$\left[\alpha_{1},\alpha_{2},\dots,\alpha_{T/\tau}\right],\quad\alpha_{i}\in\mathbb{R},$ (4)
[…]
The above equation ensures the final total sequence length (in KV cache memory) is $C_{max}$. Note that we do not consider memory usage for other operations since for long sequence inference the KV cache occupies the most GPU memory Hooper et al. (2024).
We employ chunk-based processing instead of single-frame processing to enhance the robustness of the allocation process and reduce memory overhead in temporal-adaptive allocation, as detailed in Section 4.4.
Due to the autoregressive nature of LLMs, chunked prefilling is applied to each chunk, which is functionally equivalent to standard prefilling Zeng et al. (2024b). During the $i$-th iteration, chunk $i$ is first prefilled. For each layer $l$, the query states of the prompt $\mathbf{Q}^{(l)}_{i}\in\mathbb{R}^{L\times d}$ and the KV caches of chunk $i$ $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}\in\mathbb{R}^{h\times\tau N\times d}$ are stored, where $h$ is the number of heads. These, along with the chunk’s compression ratio $\alpha_{i}$, are processed by the layer-adaptive allocation module to determine the compression ratio for each layer:
$\left[\alpha_{i}^{(1)},\alpha_{i}^{(2)},\dots,\alpha_{i}^{(l)}\right],\quad\alpha_{i}^{(l)}\in\mathbb{R},$ (6)
s.t. $\frac{\alpha_{i}^{(1)}+\alpha_{i}^{(2)}+\dots+\alpha_{i}^{(l)}}{L}=\alpha_{i}.$ (7)
Finally, token compression is applied to the visual KV caches of chunk $i$, deriving the compressed KV cache $\hat{\mathbf{K}}^{(l)}_{i},\hat{\mathbf{V}}^{(l)}_{i}\in\mathbb{R}^{\alpha_{i}^{(l)}\tau N\times d}$. The prompt states are dropped except in the last chunk.
4.3 Temporal-adaptive Allocation
Given chunked video frames $\mathcal{M}$ and maximal context length $C_{max}$, this module calculates the compression ratio for each chunk.
For video features of the $i$-th chunk $\mathbf{M}_{i}\in\mathbb{R}^{\tau\times N\times d}$, we first calculate the distance between adjacent frames $\mathbf{d}_{i}\in\mathbb{R}^{\tau-1}$:
$\mathbf{d}_{i}[t]=1-\sum_{j=1}^{N}\frac{\textmd{Sim}\left(\mathbf{M}_{i}[t,j],\mathbf{M}_{i}[t+1,j]\right)}{N},$ (8)
where $\textmd{Sim}(\cdot)$ represents the cosine similarity. We then average $\mathbf{d}_{i}$ among its $\tau-1$ frames to get the averaged distance of $i$-th chunk $\bar{d_{i}}\in\mathbb{R}$, which reflects the temporal redundancy within the chunk. Finally, the compression ratio $\alpha_{i}$ for each chunk is computed by allocating the maximal context length $C_{\text{max}}$ proportionally to the mean distances:
$\alpha_{i}=\frac{C_{max}-L}{TN}\cdot\frac{\bar{d_{i}}}{\sum_{i=1}^{T/\tau}\bar{d_{i}}}.$ (9)
4.4 Layer-adaptive Allocation
When prefilling chunk $i$ in the $l$-th LLM layer, we store the query states of the prompt $\mathbf{Q}^{(l)}_{i}$, KV cache of chunk $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}$. This module calculates the compression ratio for chunk $i$ in each layer.
In the $l$-th layer, we first calculate the attention score between prompt and the video tokens $\textbf{A}_{i}^{(l)}\in\mathbb{R}^{h\times L\times\tau N}$. We then calculate the head-averaged accumulated scores along all prompt tokens to measure the significance score of each token to the prompt, $\mathbf{a}_{i}^{(l)}\in\mathbb{R}^{\tau N}$:
$\mathbf{a}_{i}^{(l)}=\sum_{j=1}^{L}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}_{i}^{(l)}[i,j]}.$ (10)
To measure the significance of each layer, we calculate the number of tokens with large significance scores, denoted as $s_{i}^{(l)}\in\mathbb{Z}$:
$s_{i}^{(l)}$ $=\sum_{j=1}^{\tau N}\mathbb{1}\left(\mathbf{a}_{i}^{(l)}[j]>\hat{a}_{i}\right),$ (11)
[…]
$K$ $=\alpha_{i}\tau NL.$ (13)
where $\mathbb{1}(\cdot)\in\{0,1\}$ is the indicator function, $\mathop{\mathrm{K\text{-}th}}(\cdot)$ denotes the $K$-th largest value in the vector, and ||||| | denotes vector concatenation operation. Finally, we allocate the compression ratio of each layer by re-weighting the total compression ratio of current $\alpha_{i}$ in each layer:
$\alpha_{i}^{(l)}$ $=\frac{w_{i}^{(l)}}{\sum_{k=1}^{l}w_{i}^{(k)}}\alpha_{i}L,$ (14)
$w_{i}^{(l)}$ $=\frac{s_{i}^{(l)}}{\sum_{k=1}^{l}s_{i}^{(k)}}.$ (15)
Note that sometimes the $\hat{w}_{i}^{(k)}$ above might be too small. To ensure numerical stability, we introduce a minimal weight $\epsilon=0.01$ and compute the re-normalized re-weighting factor $\hat{w}_{i}^{(l)}$:
$\hat{w}_{i}^{(l)}=\frac{\max(w_{i}^{(l)}-\epsilon,0)}{\sum_{k=1}^{l}\max(w_{i}^{(k)}-\epsilon,0)}(1-L\epsilon)+\epsilon.$ (16)
For memory-efficient implementation, we calculate Eqn. (10) after each layer.
4.5 Token Compression
After prefilling the $i$-th chunk, we first drop the prompt tokens in the KV cache (except the last chunk). Based on the compression ratio derived from Eqn. 14, we then compress video tokens by selecting tokens with the top significant scores and then update the KV cache in each layer $\mathbf{K}^{(l)}$, $\mathbf{V}^{(l)}$:
$\mathcal{I}$ $=\text{ArgTopK}(\mathbf{a}_{i}^{(l)}),\quad K=\alpha_{i}^{(l)}\tau N,$ (17)
$\mathbf{K}^{(l)}$ $\leftarrow\left[\mathbf{K}^{(l)}\parallel\mathbf{K}_{i}^{(l)}[:,\mathcal{I}]\right],$ (18)
$\mathbf{V}^{(l)}$ $\leftarrow\left[\mathbf{V}^{(l)}\parallel\mathbf{V}_{i}^{(l)}[:,\mathcal{I}]\right].$ (19)
where $\text{ArgTopK}(\cdot)$ denotes the indices of $K$ elements with the largest value in the vector.
We also provide a theoretical guarantee for our layer-wise budget allocation method. See Appendix A for more details.
Theorem 4.1.
Let $I_{i}^{(l)}\in\{0,1\}$ denotes whether token $i$ in layer $l$ is kept during compression. Given the token sequence budget $\sum_{l}\sum_{i}I_{i}^{(l)}=K$, making token compression choices $\left\{\mathbf{I}_{*}^{(l)}\right\}_{l=1}^{L}$ based on top $K$ values in $\left\{A_{i}^{(l)}\right\}$ can achieve a near-optimal minimization of the upper bound of token compression loss to $\epsilon_{*}^{(l)}$:
$\epsilon_{*}^{(l)}\leq 2C+2C\left(\frac{\epsilon_{opt}^{(l)}}{2C}-1\right)^{1-\frac{1}{e}},$ (20)
where $\epsilon_{opt}^{(l)}$ is the theoretical minimal of $\epsilon^{(l)}$ and $C$ is a constant.
5 Experiments
VideoMME MLVU LongVideoBench LVBench
Model LLM Size Long Overall dev val val
GLM-4V-Plus - - 70.8 - - 58.7
GPT-4o - 65.3 71.9 64.6 66.7 27.0
Gemini-1.5-Pro - 67.4 75.0 - 64.0 33.1
VITA-1.5 7B 47.1 56.1 - - -
mPLUG-Owl3 7B 50.1 59.3 63.7 52.1 -
NVILA 8B 54.8 64.2 70.1 57.7 -
ByteVideoLLM 14B 56.4 64.6 70.1 - -
TPO 7B 55.4 65.6 71.1 60.1 -
VideoLLaMA3 7B - 66.2 73.0 59.8 45.3
LLaVA-Video 7B 52.4 63.3 67.0 58.2 43.1
LLaVA-Video+AdaR E T A K E 7B 53.9 64.0 70.6 59.6 49.6
Qwen2-VL 7B 53.8 63.3 66.9 55.6 42.4
QWen2-VL+AdaR E T A K E 7B 56.4 64.2 72.0 57.2 48.9
Qwen2.5-VL 7B 55.6 65.4 70.2 59.5 45.3
QWen2.5-VL+AdaR E T A K E 7B 58.3 67.7 75.0 62.6 51.2
LLaVA-OneVision 72B 60.0 66.3 68.0 61.3 -
Oryx-1.5 32B 59.3 67.3 72.3 62.0 30.4
Aria 8x3.5B 58.8 67.6 70.6 65.3 -
LLaVA-Video 72B 61.5 70.6 74.4 61.9 -
Qwen2-VL 72B 62.2 71.2 - 60.4 41.3
InternVL2.5 72B 62.6 72.1 75.7 63.6 43.6
Qwen2.5-VL 72B 63.9 72.6 74.6 65.9 47.3
Qwen2.5-VL+AdaR E T A K E 72B 65.0 73.5 78.1 67.0 53.3
Table 1: Performance comparison on long video understanding. AdaR
E
T
A
K
E
achieves consistent gains when integrated into various MLLMs.
5.1 Benchmarks and Implementations
Video-MME. Video Multi-Modal Evaluation Fu et al. (2024) is a pioneering benchmark designed for evaluating video analysis, with diverse video types, and durations. It comprises 900 videos totaling 256 hours, with 2,700 manually labeled complex multiple-choice question-answer pairs across 30 subfields. It has three subsets of different durations: short ($<$ 2min), medium (4min $\sim$ 15min), and long (30min $\sim$ 60min). MLVU. Multi-task Long Video Understanding Benchmark (MLVU) Zhou et al. (2024) has the widest range of video length ranging from 3 minutes to 2 hours. MLVU includes nine evaluation tasks including topic reasoning, anomaly recognition, video summarization, and plot question-answering. LongVideoBench Wu et al. (2024) is a benchmark for long-context video understanding, consisting of videos up to one hour in length. It includes 3,763 videos with 6,678 annotated multiple-choice questions across 17 categories, focusing on referring reasoning that requires retrieving and analyzing detailed multimodal information from specific temporal segments. LVBench. LVBench Wang et al. (2024b) is a comprehensive benchmark for long video understanding, with an average video length of 4,101 seconds—4 times longer than VideoMME Fu et al. (2024) and 5 times longer than MLVU Zhou et al. (2024). It includes 1,549 annotated multiple-choice question-answer pairs covering a wide range of tasks, including entity recognition, event understanding, key information retrieval, temporal grounding, and reasoning.
Implementation Details. We integrated AdaR E T A K E into various MLLMs, including LLaVA-Video-7B Zhang et al. (2024e), QWen2VL-7B Wang et al. (2024a), QWen2.5VL-7B, and QWen2.5VL-72B. We densely sampled the video at 2 frames per second (fps), with a maximum of 2048 and 1024 frames for 7B and 72B models, respectively. For our main results (Section 5.2), we chose the maximal context length $C_{max}$ as 16K. In the ablation studies (Section5.3), we reduced the maximum number of sampled frames to 1024 and the context length to 1K without specification. The evaluation is conducted using LMMs-Eval Zhang et al. (2024b). All experiments were conducted on 8 Ascend 910B NPUs.
5.2 Main Results
VideoMME MLVU LVBench
Method Long Overall val val
FastV 53.5 61.2 63.2 42.3
FitPrune 53.6 61.2 63.6 42.0
LOOK-M 53.6 61.0 63.8 42.6
SparseVLM 54.4 60.7 63.0 43.9
PyramidDrop 53.1 60.5 63.7 41.6
VL-Cache 53.2 61.3 64.5 42.4
AdaR
E
T
A
K
E
55.1 62.2 65.6 44.8
Table 2: Comparison with other token compression methods for MLLMs. AdaR
E
T
A
K
E
outperforms existing approaches by employing a theoretically grounded budget distribution mechanism, in contrast to heuristic or suboptimal allocation strategies.
Model Max frames Context length VideoMME-L MLVU LVBench $\Delta_{avg}$
0 QWen2VL-7B 128 9K 51.2 63.5 40.1 -
1 +token compression 128 1K 50.6 62.7 39.2 -1.1
2 +scale up frames 1024 1K 53.8 63.9 42.3 +2.5
3 +layer-wise allocation 1024 1K 54.3 64.6 43.5 +0.8
4 +temporal allocation 1024 1K 55.1 65.6 44.8 +1.1
5 +scale up context length 1024 16K 56.0 71.7 48.0 +3.6
Table 3: Ablation study on different components in our method. Token compression enables richer information capture, optimized compression allocation improves efficiency, and extended context length significantly enhances performance.
MLVU LVBench
Method Max Frames Max Context Length NQA AO KIR TG
LLaVA-Video-7B 128 25K 74.2 55.6 37.5 36.8
LLaVA-Video-7B+AdaR E T A K E 1024 16K 75.1 60.6 51.2 43.2
Qwen2-VL-7B 256 18K 81.9 49.0 44.3 40.5
QWen2-VL-7B+AdaR E T A K E 1024 16K 82.7 60.2 52.9 42.7
Table 4: Ablation studies on MLVU and LVBench datasets, evaluating fine-grained perception capabilities across Needle QA (NQA), Action Order (AO), Action Count (AC), Key Information Retrieval (KIR), and Temporal Grounding (TG).
Comparision with SoTAs. We integrated AdaR E T A K E with various MLLMs and compared their results with existing long video understanding methods in Table 1. The average improvements on the VideoMME, MLVU, and LVBench datasets are 1.2%, 2.8%, and 6.2%, respectively, with the most significant gains on LVBench. Given that LVBench has the longest average video duration (5x that of MLVU), we hypothesize that our method’s ability to effectively compress visual tokens enables MLLMs to process longer and more informative visual sequences, leading to greater improvements with longer video content.
Generalization for various MLLMs. When integrated with different MLLMs of different sizes, AdaR E T A K E brings consistent improvements, demonstrating its generality. With the help of AdaR E T A K E , both the 7B and 72B variants of QWen2.5-VL achieve state-of-the-art results within their respective model sizes. The 7B model sees an average improvement of 2.3%, while the 72B model achieves a 1.5% gain, demonstrating the scaling ability of our method into larger size.
Comparison with other token compression methods. As shown in Table 2, AdaR E T A K E demonstrates distinct advantages over existing MLLM token compression approaches.
Baseline methods FastV and FitPrune employ accumulated attention scores to evict tokens, while SparseVLM enhances this paradigm through partial token recycling.
PyramidDrop, VL-Cache, and our method address compression ratio allocation. However, PyramidDrop’s layer-wise monotonic budget allocation contradicts our layer importance observations in Section 3, leading to suboptimal performance. While VL-Cache improves through heuristic-based dynamic allocation, our method is theoretically grounded, achieving superior results.
5.3 Ablation Studies
Ablation studies on temporal and layer-wise adaptive allocation. To identify the sources of performance improvements in our model, we conducted ablation studies, as summarized in Table 3. In the table, #0 represents the baseline model. In #1, we directly incorporates token compression into baseline model, and in #2, we increase the number of sampled frames while keeping the maximum context length fixed. In #3 and #4 we applly varying compression ratios across different layers and different frames respectively. Finally, #5 extends the context length. First, comparing rows 0,1 and 1,2 reveals that token compression introduces a slight performance drop (-1.1% on average). However, it enables the model to process more frames within the same context length, capturing richer information and ultimately yielding a net performance gain (2.5% on average versus -1.1%). 2) Comparing rows 2,3 and 3,4 shows that our strategy of distributing the compression ratio across frames and layers enhances performance (by 1.1% and 0.8% on average, respectively), confirming the effectiveness of our AdaR E T A K E . 3) Comparing rows 4 and 5 demonstrates that scaling the context length to the typical upper limit of MLLMs Shen et al. (2024) further improves performance significantly, with an average gain of 3.6%.
Perception ability on temporal details. To assess the effectiveness of token compression algorithms in preserving critical temporal details, we conducted ablation studies on the MLVU and LVBench datasets, focusing on Needle QA, Action Order, Key Information Retrieval, and Temporal Grounding. We compared baseline models LLaVA-Video-7B and QWen2-VL-7B, maximizing frame sampling within their constraints (128 and 256 frames, respectively). Results are shown in Table 4. Our analysis reveals three key findings: 1) Despite token compression via AdaR E T A K E , increasing the maximum sampled frames improved grounding abilities (Needle QA and Temporal Grounding) without compromising temporal order perception (Action Order). This indicates that AdaR E T A K E enhances model performance while strengthening fine-grained temporal capabilities. 2) The improvement in MLVU’s Action Order category was significantly higher than in Needle QA (8% vs. 0.8% on average). We attribute this to our method’s ability to sample more frames through token compression, thus a denser frame sampling is enabled, which greatly enhances action understanding Li et al. (2024b). 3) In LVBench, under similar baselines, Key Information Retrieval demonstrated a significantly higher improvement compared to Temporal Grounding, with average gains of 11.2% versus 4.3%. We hypothesize that token compression enhances information density, which strengthens comprehensive understanding. We believe this can explain why Key Information Retrieval, a task requiring deeper comprehension, benefits more than perceptual tasks like Temporal Grounding in our results.
6 Conclusion
We introduce AdaR E T A K E , a training-free method for adaptive redundancy reduction in MLLMs. By dynamically allocating compression ratios across frames and model layers, AdaR E T A K E achieves more efficient video token compression. This allows us to scale up to more frames and extract valuable information within the same computational budget. Integrated into state-of-the-art MLLMs, AdaR E T A K E enables processing of up to 2,048 frames and outperforms existing methods on benchmarks such as VideoMME, MLVU, LongVideoBench, and LVBench by a large margin.
7 Limitations
While AdaR E T A K E can be integrated into most MLLMs, it may also inherit their inherent limitations, such as factual inaccuracies, biases, and hallucinations.
Appendix A Proof of Token Compression Loss
A.1 Preliminaries
LLMs are characterized by an autoregressive generation mode, where each step involves using the last token to predict the next token. For a formal representation of our approach, we denote $\mathbf{K}^{(l)}$,$\mathbf{V}^{(l)}\in\mathbb{R}^{n-1\times d},i=1,2,...,L$ as the video KV cache in and $\mathbf{x}^{(l)}\in\mathbb{R}^{d}$ as the last token used as input at the current time step, where $n-1$ is the prefilled sequence length, $L$ is the number of layers and $d$ is the hidden state dimension. Note that we ignore the number of attention heads for simplification.
During attention forward in the $l$-th layer, the input token is first mapped into its query, key, and value states:
$\mathbf{q}^{(l)}$ $=\mathbf{x}^{(l)}\mathbf{W}_{Q}^{(l)},$ (21)
$\mathbf{k}^{(l)}$ $=\mathbf{x}^{(l)}\mathbf{W}_{K}^{(l)},$ (22)
$\mathbf{v}^{(l)}$ $=\mathbf{x}^{(l)}\mathbf{W}_{V}^{(l)},$ (23)
where $\mathbf{W}_{Q}^{(l)},\mathbf{W}_{K}^{(l)},\mathbf{W}_{V}^{(l)}$ are transformation matrices. Then, the previous KV Cache is updated:
$\mathbf{K}^{(l)}$ $\leftarrow\left[\mathbf{K}^{(l)}||\mathbf{k}^{(l)}\right],$ (24)
$\mathbf{V}^{(l)}$ $\leftarrow\left[\mathbf{V}^{(l)}||\mathbf{v}^{(l)}\right],$ (25)
where ||||| | denotes vector concatenation. Finally, the output of the $l$-th layer $\mathbf{y}^{(l)}\in\mathbb{R}^{d}$ is computed:
$\mathbf{y}^{(l)}$ $=\mathbf{A}^{(l)}\mathbf{V}^{(l)}\mathbf{O}_{V}^{(l)},$ (26)
$\mathbf{A}^{(l)}$ $=\textmd{Softmax}\left(\mathbf{q}^{(l)}(\mathbf{K}^{(l)})^{T}\right),$ (27)
where $\mathbf{O}_{V}^{(l)}$ is the transformation matrics, $\mathbf{A}^{(l)}$ is the attention weights.
A.2 Attention Output after Compression
We define token compression choice $\mathbf{I}^{(l)}\in\{0,1\}^{n}$ to represent which token to preserve during compression, where the value of its $i$-th element satisfies:
$I_{i}^{(l)}=\begin{cases}1,&\text{if }K_{i}^{(l)}\text{ and }V_{i}^{(l)}\text{are retained},\\ 0,&\text{if }K_{i}^{(l)}\text{ and }V_{i}^{(l)}\text{ are dropped},\end{cases}$ (28)
where $K_{i}^{(l)}$ and $V_{i}^{(l)}$ is the $i$-th element in $\textbf{K}^{(l)}$ and $\textbf{V}^{(l)}$, respectively. Based on this, the output of the $l$-th layer after compression $\hat{\mathbf{y}}^{(l)}\in\mathbb{R}^{d}$ can be represented in a concise format according to the lemma below. Note that this lemma borrows insights from Theorem 1 in AdaKV Feng et al. (2024), we extend it to the multi-layer scenario.
Lemma A.1.
Given token compression choice $\mathbf{I}^{(l)}$, the output of the $l$-th layer after compression $\hat{y}^{(l)}$ can be rewritten as:
$\hat{\mathbf{y}}^{(l)}=\frac{\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}}{||\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}||_{1}}\textbf{V}^{(l)}\textbf{W}_{O}^{(l)},$ (29)
where $\hat{\textbf{A}}^{(l)}=\hat{\mathbf{q}}^{(l)}(\mathbf{K}^{(l)})^{T}$, $\hat{\mathbf{q}}^{(l)}$ is the input of the $l$-th attention layer after KV cache in the $(l-1)$-th layer is compressed.
Proof.
Considering that Softmax function is:
$\text{Softmax}(x)_{j}=\frac{\exp(x_{j})}{\sum_{j}\exp(x+j)},$ (30)
we can use $\infty$- ∞ to represent a token that is dropped during softmax calculation. Thus, the attention weight in the $l$-th layer after token compression is:
$\tilde{\textbf{A}}^{(l)}=\text{Softmax}\left(-\infty\odot(\mathbf{1}-\mathbf{I}^{(l)})+\hat{\textbf{s}}^{(l)}\right),$ (31)
where $\hat{\textbf{s}}^{(l)}$ is the attention logits before compression:
$\hat{\textbf{s}}^{(l)}=\hat{\mathbf{q}}^{(l)}(\mathbf{K}^{(l)})^{T}.$ (32)
The $i$-th element of $\tilde{\textbf{A}}^{(l)}$ is:
$\tilde{A}_{i}^{(l)}$ $=\frac{\exp(\hat{s}_{i}^{(l)}-\infty\odot(1-I_{i}^{(l)}))}{\sum_{j}\exp(\hat{s}_{j}^{(l)}-\infty\odot(1-I_{j}^{(l)}))},$ (33)
$=\frac{I_{i}^{(l)}\exp(\hat{s}_{i}^{(l)})}{\sum_{j}I_{j}^{(l)}\exp(\hat{s}_{j}^{(l)})},$ (34)
$=\frac{I_{i}^{(l)}\exp(\hat{s}_{i}^{(l)})}{\sum_{j}\exp(\hat{s}_{j}^{(l)})}\frac{\sum_{j}\exp(\hat{s}_{j}^{(l)})}{\sum_{j}I_{j}^{(l)}\exp(\hat{s}_{j}^{(l)})},$ (35)
where $\hat{s}_{i}^{(l)}$ is the $i$-th element in $\hat{\textbf{s}}^{(l)}$.
Denote $\hat{\textbf{A}}^{(l)}=\hat{\mathbf{q}}^{(l)}(\mathbf{K}^{(l)})^{T}$, we can get further simplify $\tilde{A}_{i}^{(l)}$ above:
$\tilde{A}_{i}^{(l)}$ $=I_{i}^{(l)}\hat{A}_{i}^{(l)}\frac{\sum_{j}\exp(\hat{s}_{j}^{(l)})}{\sum_{j}I_{j}^{(l)}\exp(\hat{s}_{j}^{(l)})},$ (36)
$=\frac{I_{i}^{(l)}\hat{A}_{i}^{(l)}}{||\hat{\textbf{A}}^{(l)}\odot\textbf{I}^{(l)}||_{1}},$ (37)
where $\hat{A}_{i}^{(l)}$ is the $i$-th element in $\hat{\textbf{A}}^{(l)}$.
Then we can obtain the simplified form of the attention weight after token compression in Eqn. (31):
$\tilde{\textbf{A}}^{(l)}=\frac{\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}}{||\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}||_{1}}.$ (38)
Thus:
$\hat{\textbf{y}}^{(l)}=\frac{\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}}{||\hat{\textbf{A}}^{(l)}\odot\mathbf{I}^{(l)}||_{1}}\textbf{V}^{(l)}\textbf{W}_{O}^{(l)}.$ (39)
A.3 Upper Bound of Token Compression Loss
We first study the compression error at each layer and then the compression error at the last layer, i.e., the token compression loss.
To measure the information loss during token compression, we study layer compression error, i.e., the $L_{1}$ distance of the output in the $l$-th layer before and after compression $\mathcal{D}^{(l)}$:
$\mathcal{D}^{(l)}=||\mathbf{y}^{(l)}-\hat{\mathbf{y}}^{(l)}||_{1}.$ (40)
The layer compression error in the 1st layer can be bounded by the following lemma:
Lemma A.2.
Given token compression choice $\mathbf{I}^{(1)}$, the 1st layer compression error $\mathcal{D}^{(1)}$ can be bounded by:
$\mathcal{D}^{(1)}\leq 2C^{(1)}-2C^{(1)}\sum_{i=1}^{n}I_{i}^{(1)}A_{i}^{(1)},$ (41)
where $C^{(1)}=\|\textbf{V}^{(1)}\textbf{W}_{O}^{(1)}\|_{\infty}$ is a constant.
Proof.
By expanding 1st layer compression error Eqn. (40) with Lemma A.1 and using $\hat{\mathbf{A}}^{(1)}=\mathbf{A}^{(1)}$, we can obtain:
$\mathcal{D}^{(1)}=\\ \left\|(\mathbf{1}-\frac{\mathbf{I}^{(1)}}{||\mathbf{A}^{(1)}\odot\mathbf{I}^{(1)}||_{1}})\odot\mathbf{A}^{(1)}\textbf{V}^{(1)}\textbf{W}_{O}^{(1)}\right\|_{1},$ (42)
$\leq\left\|(\mathbf{1}-\frac{\mathbf{I}^{(1)}}{||\mathbf{A}^{(1)}\odot\mathbf{I}^{(1)}||_{1}})\odot\textbf{A}^{(l)}\right\|_{1}\left\|\textbf{V}^{(1)}\textbf{W}_{O}^{(1)}\right\|_{\infty},$ (43)
$\leq C^{(1)}\left\|(\mathbf{1}-\frac{\mathbf{I}^{(1)}}{||\mathbf{A}^{(1)}\odot\mathbf{I}^{(1)}||_{1}})\odot\textbf{A}^{(l)}\right\|_{1}.$ (44)
Denote $C^{(1)}=||\textbf{V}^{(l)}\textbf{W}_{O}^{(l)}||_{\infty}$, and we apply Hölder’s inequality to derive the first inequality.
Denote $F^{(1)}=||\mathbf{A}^{(1)}\odot\mathbf{I}^{(1)}||_{1}\in(0,1]$, and consider the definition of $I_{i}^{(l)}$ in Eqn. (28), we can further simplify the compression error by expanding $\mathbf{A}^{(1)}$:
$\mathcal{D}^{(1)}\leq C^{(1)}\left\|(\mathbf{1}-\frac{\mathbf{I}^{(1)}}{||\mathbf{A}^{(1)}\odot\mathbf{I}^{(1)}||_{1}})\odot\textbf{A}^{(l)}\right\|_{1},$ (45)
$=C^{(1)}\sum_{i=1}^{n}\frac{|F^{(1)}-I_{i}^{(1)}|A_{i}^{(1)}}{F^{(1)}},$ (46)
$=C^{(1)}\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=0\end{subarray}}^{n}A_{i}^{(1)}+C^{(1)}\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=1\end{subarray}}^{n}\frac{(1-F^{(1)})A_{i}^{(1)}}{F^{(1)}},$ (47)
$=C^{(1)}\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=0\end{subarray}}^{n}A_{i}^{(1)}+C^{(1)}\left(\frac{\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=1\end{subarray}}^{n}A_{i}^{(1)}}{F^{(1)}}-\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=1\end{subarray}}^{n}A_{i}^{(1)}\right).$ (48)
Considering that there exists the following relationship between $I_{i}^{(1)}$ and $A_{i}^{(1)}$:
$F^{(1)}=\sum_{i=1}^{n}A_{i}^{(1)}I_{i}^{(1)}=\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=1\end{subarray}}^{n}A_{i}^{(1)}.$ (49)
The compression error can be further simplified:
$\mathcal{D}^{(1)}\leq C^{(1)}\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=0\end{subarray}}^{n}A_{i}^{(1)}+C^{(1)}\left(1-\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=1\end{subarray}}^{n}A_{i}^{(1)}\right),$ (50)
$=2C^{(1)}\sum_{\begin{subarray}{c}i=1\\ I_{i}^{(1)}=0\end{subarray}}^{n}A_{i}^{(1)},$ (51)
$=2C^{(1)}\sum_{i=1}^{n}(1-I_{i}^{(1)})A_{i}^{(1)},$ (52)
$=2C^{(1)}-2C^{(1)}\sum_{j=1}^{n}I_{i}^{(1)}A_{i}^{(1)}.$ (53)
To figure out how layer compression error propagates through layers, we have the following lemma:
Lemma A.3.
Let $C^{(l)}=\|\textbf{V}^{(l)}\textbf{W}_{O}^{(l)}\|_{\infty}$ and assume $14C^{(l)}>$ 1 for $l=1,2,\dots,L$. In the $l$-th layer, given token compression choice $\mathbf{I}^{(l)}$, the layer compression error $\mathcal{D}^{(l)}$ is bounded by $\epsilon^{(l)}$:
$\mathcal{D}^{(l)}\leq\epsilon^{(l)}=2C^{(l)}-2C^{(l)}\prod_{k=1}^{(l)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}.$ (54)
Proof.
Following the same logic with proof of Lemma A.2, the $l$-th layer compression error is bounded by:
$\mathcal{D}^{(l)}\leq 2C^{(l)}-2C^{(l)}\sum_{i=1}^{n}I_{i}^{(l)}\hat{A}_{i}^{(l)}.$ (55)
Let $\delta\mathbf{q}^{(l)}\in\mathbb{R}^{n}$ be the perturbation of the input of the $l$-th attention layer after video compression:
$\hat{\mathbf{q}}^{(l)}=(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}).$ (56)
Considering that $\hat{\mathbf{q}}^{(l)}$ is obtained by applying an MLP followed by a linear layer to $\hat{\mathbf{y}}^{(l-1)}$ and $||\hat{\mathbf{y}}^{(l-1)}-\mathbf{y}^{(l-1)}||_{1}\leq\epsilon^{(l-1)}$, the perturbation $\delta\mathbf{q}^{(l)}$ is also bound by:
$||\delta\mathbf{q}^{(l)}||_{1}\leq\epsilon^{(l-1)}.$ (57)
Then, $\hat{\textbf{A}}^{(l)}$ can be calculated with:
$\hat{\textbf{A}}^{(l)}=\hat{\mathbf{q}}^{(l)}(\mathbf{K}^{(l)})^{T}=(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)})(\mathbf{K}^{(l)})^{T}.$ (58)
The $i$-th element in $\hat{\textbf{A}}^{(l)}$ is:
$\hat{A}_{i}^{(l)}=\frac{\exp\left[\left(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}\right)\left(\mathbf{K}_{i}^{(l)}\right)^{T}\right]}{\sum_{j=1}^{n}\exp\left[\left(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}\right)\left(\mathbf{K}_{j}^{(l)}\right)^{T}\right]},$ (59)
where $\mathbf{K}_{i}^{(l)}$ is the $i$-th row vector in $\mathbf{K}^{(l)}$.
Using the Hölder’s inequality, we can bound the perturbation term:
$\left|\delta\mathbf{q}^{(l)}\left(\mathbf{K}_{j}^{(l)}\right)^{T}\right|\leq||\delta\mathbf{q}^{(l)}||_{1}||\mathbf{K}_{j}^{(l)}||_{\infty}.$ (60)
Based on Eqn. (57), we obtain:
[…]
And the exponent in Eqn. (59) satisfies:
$\mathbf{q}^{(l)}\left(\mathbf{K}_{j}^{(l)}\right)^{T}-\epsilon^{(l-1)}||\mathbf{K}_{j}^{(l)}||_{\infty}\leq$ (62)
$\left(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}\right)\left(\mathbf{K}_{j}^{(l)}\right)^{T}\leq$
$\mathbf{q}^{(l)}\left(\mathbf{K}_{j}^{(l)}\right)^{T}+\epsilon^{(l-1)}||\mathbf{K}_{j}^{(l)}||_{\infty}.$
Therefore, the $i$-th element in Eqn. (59) is bounded by:
$\hat{A}_{i}^{(l)}=\frac{\exp\left[\left(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}\right)\left(\mathbf{K}_{i}^{(l)}\right)^{T}\right]}{\sum_{j=1}^{n}\exp\left[\left(\mathbf{q}^{(l)}+\delta\mathbf{q}^{(l)}\right)\left(\mathbf{K}_{j}^{(l)}\right)^{T}\right]},$ (63)
$\geq\frac{\exp\left[\mathbf{q}^{(l)}\left(\mathbf{K}_{i}^{(l)}\right)^{T}-\epsilon^{(l-1)}||\mathbf{K}_{i}^{(l)}||_{\infty}\right]}{\sum_{j=1}^{n}\exp\left[\mathbf{q}^{(l)}\left(\mathbf{K}_{j}^{(l)}\right)^{T}+\epsilon^{(l-1)}||\mathbf{K}_{j}^{(l)}||_{\infty}\right]},$ (64)
$\geq A_{i}^{(l)}\frac{\exp\left[-\epsilon^{(l-1)}\max\left\{||\mathbf{K}_{i}^{(l)}||_{\infty}\right\}\right]}{\exp\left[\epsilon^{(l-1)}\max\left\{||\mathbf{K}_{i}^{(l)}||_{\infty}\right\}\right]},$ (65)
$=A_{i}^{(l)}\exp\left[-2\epsilon^{(l-1)}\max\left\{||\mathbf{K}_{i}^{(l)}||_{\infty}\right\}\right],$ (66)
$\geq A_{i}^{(l)}\exp\left[-2\epsilon^{(l-1)}\right],$ (67)
$\geq A_{i}^{(l)}\left[-2\epsilon^{(l-1)}+1\right].$ (68)
We do not know the exact form of $\epsilon^{(l-1)}$ by now. But we can use mathematical induction to prove it. The base case $l-1=1$ has been proved in Lemma A.2. We hypothesize that the following inductive hypothesis holds:
$\epsilon^{(l-1)}=2C^{(l-1)}-2C^{(l-1)}\prod_{k=1}^{(l-1)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}.$ (69)
Combine Eqn. (55), Eqn. (68), and assumption $14C^{(l-1)}>$ 1, we can prove the inductive step:
$\mathcal{D}^{(l)}\leq 2C^{(l)}-2C^{(l)}\sum_{i=1}^{n}I_{i}^{(l)}\hat{A}_{i}^{(l)},$ (70)
$\leq 2C^{(l)}-2C^{(l)}\left[-2\epsilon^{(l-1)}+1\right]\sum_{i=1}^{n}I_{i}^{(l)}A_{i}^{(l)},$ (71)
$=2C^{(l)}-$ (72)
$2C^{(l)}\left[-4C^{(l-1)}+4C^{(l-1)}\prod_{k=1}^{(l-1)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}+1\right]$
$\sum_{i=1}^{n}I_{i}^{(l)}A_{i}^{(l)},$
$\leq 2C^{(l)}-2C^{(l)}\left[\prod_{k=1}^{(l-1)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}\right]\sum_{i=1}^{n}I_{i}^{(l)}A_{i}^{(l)},$ (73)
$=2C^{(l)}-2C^{(l)}\left[\prod_{k=1}^{(l)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}\right].$ (74)
Finally,
$\mathcal{D}^{(l)}\leq 2C^{(l)}-2C^{(l)}\prod_{k=1}^{(l)}\sum_{i=1}^{n}I_{i}^{(k)}A_{i}^{(k)}=\epsilon^{(l)}.$ (75)
By the principle of mathematical induction, since the statement is true for the base case and the inductive step ensures it holds for all subsequent values, the statement is proven for all $l$ in the domain. ∎
To measure the final information loss during token compression, we define the compression loss $\mathcal{L}$ as the $L_{1}$ distance of the output in the last (i.e., $L$-th) layer before and after compression:
$\mathcal{L}=||\mathbf{y}^{(l)}-\hat{\mathbf{y}}^{(l)}||_{1}.$ (76)
By applying Lemma A.3 for $l=L$. The following theorem can characterize the upper bound of the compression loss.
Theorem A.4.
Let $C^{(l)}=\|\textbf{V}^{(l)}\textbf{W}_{O}^{(l)}\|_{\infty}$ and assume $14C^{(l)}>$ 1 for $l=1,2,\dots,L$. Given token compression choices $\left\{\mathbf{I}^{(l)}\right\}_{l=1}^{(L)}$, the compression loss $\mathcal{L}$ can be bounded by $\epsilon^{L}$:
$\mathcal{L}\leq\epsilon^{L}=2C^{(L)}-2C^{(L)}\prod_{l=1}^{L}\sum_{i=1}^{n}I_{i}^{(l)}A_{i}^{(l)}.$ (77)
A.4 Minimizing the Upper Bound of Token Compression Loss
We can prove that by selecting tokens based on top $C_{max}$ values in $\left\{A_{i}^{l}\right\}$, we can achieve a near near-optimal minimization of the upper bound of token compression loss in Lemma A.4.
Theorem A.5.
Given the token sequence budget $\sum_{l}\sum_{i}I_{i}^{(l)}=K$, selecting token compression choices $\left\{\mathbf{I}_{*}^{(l)}\right\}_{l=1}^{(L)}$ based on top $K$ values in $\left\{A_{i}^{l}\right\}$ can achieve a near-optimal minimization of the upper bound of token compression loss to $\epsilon_{*}^{L}$:
$\epsilon_{*}^{L}\leq 2C^{(L)}+2C^{(L)}\left(\frac{\epsilon_{opt}^{L}}{2C^{(L)}}-1\right)^{1-\frac{1}{e}},$ (78)
where $\epsilon_{opt}^{L}$ is the theoretical minimal of $\epsilon^{L}$.
Proof.
To minimize $\epsilon^{L}$, we can maximize the product-sum within:
$F(\mathcal{S})=\prod_{l=1}^{L}\sum_{i\in\mathcal{S}_{l}}A^{(l)}_{i}=\exp\left(\sum_{l=1}^{L}\log\left(\sum_{i\in\mathcal{S}_{l}}A^{(l)}_{i}\right)\right),$ (79)
where $\mathcal{S}=\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{L}\}$ denotes the sets of selected elements for each layer, with $\sum_{l=1}^{L}|\mathcal{S}_{l}|=K$.
Maximizing $F(\mathcal{S})$ is equivalent to maximizing the log-product $f(\mathcal{S})$:
$f(\mathcal{S})=\sum_{l=1}^{L}\log\left(\sum_{i\in\mathcal{S}_{l}}A^{(l)}_{i}\right).$ (80)
Next, we prove its submodular property for minimization. A set function $f$ is submodular if it satisfies diminishing marginal returns:
$f(\mathcal{S}\cup\{e\})-f(\mathcal{S})\geq$ (81)
$f(\mathcal{T}\cup\{e\})-f(\mathcal{T}),\quad\forall\mathcal{S}\subseteq\mathcal{T}.$
For our problem, adding an element $e$ (from layer $l$) to $\mathcal{S}$ increases $f(\mathcal{S})$ by:
$\Delta_{f}(e\mid\mathcal{S})=\log\left(S_{l}+A^{(l)}_{e}\right)-\log(S_{l}),$ (82)
where $S_{l}=\sum_{i\in\mathcal{S}_{l}}A^{(l)}_{i}$. Since $\Delta_{f}(e\mid\mathcal{S})$ decreases as $S_{l}$ grows (diminishing returns), $f$ is monotone submodular.
[…]
$f(\mathcal{S}_{\text{greedy}})\geq\left(1-\frac{1}{e}\right)f(\mathcal{S}^{*}).$
For multiplicative objectives like $F(\mathcal{S})=e^{f(\mathcal{S})}$, this translates to:
$F(\mathcal{S}_{\text{greedy}})\geq e^{(1-1/e)\log F(\mathcal{S}^{*})}=F(\mathcal{S}^{*})^{1-1/e}.$
Therefore,
$\prod_{l=1}^{L}\sum_{i=1}^{n}{I_{*}}_{i}^{(l)}{A_{*}}_{i}^{(l)}\geq\left(\prod_{l=1}^{L}\sum_{i=1}^{n}{I_{opt}}_{i}^{(l)}{A_{opt}}_{i}^{(l)}\right)^{1-\frac{1}{e}}.$ (83)
Based on Eqn. (A.4), we can finally get:
$\epsilon_{*}^{L}\leq 2C^{(L)}+2C^{(L)}\left(\frac{\epsilon_{opt}^{L}}{2C^{(L)}}-1\right)^{1-\frac{1}{e}}.$ (84)

### v2

@@
 We propose AdaR E T A K E , a training-free method that flexibly reduces visual redundancy by allocating compression ratios among time and layers with theoretical guarantees.
-Integrated into state-of-the-art MLLMs, AdaR E T A K E improves processing capacity from 256 to 2048 frames while preserving critical information.
+AdaR E T A K E can be seamlessly integrated into existing MLLMs as a plug-and-play solution, extending their processing capacity from 256 to 2048 frames while preserving critical information.
 Experiments on VideoMME, MLVU, LongVideoBench, and LVBench datasets demonstrate that AdaR E T A K E outperforms existing methods by 2.3% and 2.8% for 7B and 72B models, respectively, with even greater improvements of 5.9% and 6.0% on the longest LVBench.
@@
 (2024b); Lin et al.
-(2024) have revolutionized video understanding.
+(2024); Wang et al.
+(2025b, a) have revolutionized video understanding.
 However, current MLLMs require hundreds of tokens to represent a single image Wang et al.
 (2024a); Li et al.
-(2024a), limiting video lengths to less than 10 minutes Shen et al.
+(2024a); Wang et al.
+(2023), limiting video lengths to less than 10 minutes Shen et al.
+(2024); Gan et al.
 (2023).
@@
 (2024d); Luo et al.
+(2024); Liu et al.
 (2018) that segment videos into shorter clips and use MLLMs with open-source tools for retrieval, aggregation, and interpretation.
@@
 4.2 General Pipeline
-Denote $T$ number of frames, $N$ number of tokens in each frame, $\tau$ number of frames in a chunk (can divide $T$), $L$ prompt length, and $C_{max}$ is a refined maximal context length.
-Given raw frames and a text prompt as input, the visual encoder and projection layer derive video features $\mathbf{M}\in\mathbb{R}^{T\times N\times d}$, and the word embedding layer derives prompt features $\mathbf{P}\in\mathbb{R}^{L\times d}$.
+Denote $T$ number of frames, $N$ number of tokens in each frame, $\tau$ number of frames in a chunk (can divide $T$), $S$ prompt length, $L$ number of LLM layers, and $C_{max}$ is a refined maximal context length.
+Given raw frames and a text prompt as input, the visual encoder and projection layer derive video features $\mathbf{M}\in\mathbb{R}^{T\times N\times d}$, and the word embedding layer derives prompt features $\mathbf{P}\in\mathbb{R}^{S\times d}$.
 We split visual features into chunks of $\tau$ frames:
@@
 During the $i$-th iteration, chunk $i$ is first prefilled.
-For each layer $l$, the query states of the prompt $\mathbf{Q}^{(l)}_{i}\in\mathbb{R}^{L\times d}$ and the KV caches of chunk $i$ $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}\in\mathbb{R}^{h\times\tau N\times d}$ are stored, where $h$ is the number of heads.
+For each layer $l$, the query states of the prompt $\mathbf{Q}^{(l)}_{i}\in\mathbb{R}^{S\times d}$ and the KV caches of chunk $i$ $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}\in\mathbb{R}^{h\times\tau N\times d}$ are stored, where $h$ is the number of heads.
 These, along with the chunk’s compression ratio $\alpha_{i}$, are processed by the layer-adaptive allocation module to determine the compression ratio for each layer:
@@
 Finally, the compression ratio $\alpha_{i}$ for each chunk is computed by allocating the maximal context length $C_{\text{max}}$ proportionally to the mean distances:
-$\alpha_{i}=\frac{C_{max}-L}{TN}\cdot\frac{\bar{d_{i}}}{\sum_{i=1}^{T/\tau}\bar{d_{i}}}.$ (9)
+$\alpha_{i}=\frac{C_{max}-S}{TN}\cdot\frac{\bar{d_{i}}}{\sum_{i=1}^{T/\tau}\bar{d_{i}}}.$ (9)
 4.4 Layer-adaptive Allocation
@@
 This module calculates the compression ratio for chunk $i$ in each layer.
-In the $l$-th layer, we first calculate the attention score between prompt and the video tokens $\textbf{A}_{i}^{(l)}\in\mathbb{R}^{h\times L\times\tau N}$.
+In the $l$-th layer, we first calculate the attention score between prompt and the video tokens $\textbf{A}_{i}^{(l)}\in\mathbb{R}^{h\times S\times\tau N}$.
 We then calculate the head-averaged accumulated scores along all prompt tokens to measure the significance score of each token to the prompt, $\mathbf{a}_{i}^{(l)}\in\mathbb{R}^{\tau N}$:
-$\mathbf{a}_{i}^{(l)}=\sum_{j=1}^{L}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}_{i}^{(l)}[i,j]}.$ (10)
+$\mathbf{a}_{i}^{(l)}=\sum_{j=1}^{S}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}_{i}^{(l)}[i,j]}.$ (10)
 To measure the significance of each layer, we calculate the number of tokens with large significance scores, denoted as $s_{i}^{(l)}\in\mathbb{Z}$:
@@
 Finally, we allocate the compression ratio of each layer by re-weighting the total compression ratio of current $\alpha_{i}$ in each layer:
-$\alpha_{i}^{(l)}$ $=\frac{w_{i}^{(l)}}{\sum_{k=1}^{l}w_{i}^{(k)}}\alpha_{i}L,$ (14)
+$\alpha_{i}^{(l)}$ $=w_{i}^{(l)}\alpha_{i},$ (14)
 $w_{i}^{(l)}$ $=\frac{s_{i}^{(l)}}{\sum_{k=1}^{l}s_{i}^{(k)}}.$ (15)
@@
 (2024b).
-All experiments were conducted on 8 Ascend 910B NPUs.
 5.2 Main Results
@@
 0 QWen2VL-7B 128 9K 51.2 63.5 40.1 -
-1 +token compression 128 1K 50.6 62.7 39.2 -1.1
+1 +token compression 128 1K 50.6 62.7 39.2 -0.8
 2 +scale up frames 1024 1K 53.8 63.9 42.3 +2.5
 3 +layer-wise allocation 1024 1K 54.3 64.6 43.5 +0.8
-4 +temporal allocation 1024 1K 55.1 65.6 44.8 +1.1
-5 +scale up context length 1024 16K 56.0 71.7 48.0 +3.6
+4 +temporal allocation 1024 1K 55.1 65.6 44.8 +1.0
+5 +scale up context length 1024 16K 56.0 71.7 48.0 +3.4
+6 +scale up frames 2048 16K 56.4 72.0 48.9 +0.6
 Table 3: Ablation study on different components in our method.
@@
 Finally, #5 extends the context length.
-First, comparing rows 0,1 and 1,2 reveals that token compression introduces a slight performance drop (-1.1% on average).
-However, it enables the model to process more frames within the same context length, capturing richer information and ultimately yielding a net performance gain (2.5% on average versus -1.1%).
-2) Comparing rows 2,3 and 3,4 shows that our strategy of distributing the compression ratio across frames and layers enhances performance (by 1.1% and 0.8% on average, respectively), confirming the effectiveness of our AdaR E T A K E .
+First, comparing rows 0,1 and 1,2 reveals that token compression introduces a slight performance drop (-0.8% on average).
+However, it enables the model to process more frames within the same context length, capturing richer information and ultimately yielding a net performance gain (2.5% on average versus -0.8%).
+2) Comparing rows 2,3 and 3,4 shows that our strategy of distributing the compression ratio across frames and layers enhances performance (by 1.0% and 0.8% on average, respectively), confirming the effectiveness of our AdaR E T A K E .
 3) Comparing rows 4 and 5 demonstrates that scaling the context length to the typical upper limit of MLLMs Shen et al.
-(2024) further improves performance significantly, with an average gain of 3.6%.
+(2024) further improves performance significantly, with an average gain of 3.4%.
 Perception ability on temporal details.
@@
 While AdaR E T A K E can be integrated into most MLLMs, it may also inherit their inherent limitations, such as factual inaccuracies, biases, and hallucinations.
+8 Acknowledgment
+This work was supported in part by the National Natural Science Foundation of China(Grant Nos.
+62376069 and 62236003), in part by the Young Elite Scientists Sponsorship Program by CAST (Grant No.
+2023QNRC001), in part by Guangdong Basic and Applied Basic Research Foundation (Grant No.
+2024A1515012027), in part by Jiangsu Science and Technology Major Program (Grant No.
+BG2024041), and in part by the Shenzhen Science and Technology Program (Grant Nos.
+KQTD20240729102207002 and ZDSYS20230626091203008).
 Appendix A Proof of Token Compression Loss
@@
 LLMs are characterized by an autoregressive generation mode, where each step involves using the last token to predict the next token.
-For a formal representation of our approach, we denote $\mathbf{K}^{(l)}$,$\mathbf{V}^{(l)}\in\mathbb{R}^{n-1\times d},i=1,2,...,L$ as the video KV cache in and $\mathbf{x}^{(l)}\in\mathbb{R}^{d}$ as the last token used as input at the current time step, where $n-1$ is the prefilled sequence length, $L$ is the number of layers and $d$ is the hidden state dimension.
+For a formal representation of our approach, we denote $\mathbf{K}^{(l)}$,$\mathbf{V}^{(l)}\in\mathbb{R}^{(n-1)\times d},i=1,2,...,L$ as the video KV cache in and $\mathbf{x}^{(l)}\in\mathbb{R}^{d}$ as the last token used as input at the current time step, where $n-1$ is the prefilled sequence length, $L$ is the number of layers and $d$ is the hidden state dimension.
 Note that we ignore the number of attention heads for simplification.
@@
 $\mathbf{v}^{(l)}$ $=\mathbf{x}^{(l)}\mathbf{W}_{V}^{(l)},$ (23)
-where $\mathbf{W}_{Q}^{(l)},\mathbf{W}_{K}^{(l)},\mathbf{W}_{V}^{(l)}$ are transformation matrices. Then, the previous KV Cache is updated:
+where $\mathbf{W}_{Q}^{(l)},\mathbf{W}_{K}^{(l)},\mathbf{W}_{V}^{(l)}\in\mathbb{R}^{d\times d}$ are transformation matrices. Then, the previous KV Cache is updated:
 $\mathbf{K}^{(l)}$ $\leftarrow\left[\mathbf{K}^{(l)}||\mathbf{k}^{(l)}\right],$ (24)